OpenStudy (anonymous):

find the coordinate of the foot of the perpendicular from the point (2,3) on the line x+y=11

OpenStudy (tamtoan):

if you are given a slope of a line, what is the slope of perpendicular line?

OpenStudy (anonymous):

-1 ?

OpenStudy (tamtoan):

ok..then you will have one equation...let's say the 2 lines intersect at point (x1,y1) ..then write equation of slope of perpendicular line... remember the point (x1,y1) is also on the original line too so you will have a system of 2 equations with 2 variables, you can solve for that

OpenStudy (anonymous):

Actually the slope of the perpendicular line is equal to -1 over the slope w/c is given. I cas her you are correct @Devanshi

OpenStudy (anonymous):

let coordinate be (x,y) x + y = 11 and (y - 3)/(x - 2) = 1 m1 m2 = -1 x = 5 y = 6

OpenStudy (tamtoan):

devanshi, what is the slope of the line x + y = 11 ?

OpenStudy (anonymous):

1

OpenStudy (tamtoan):

devanshi , slope is in the form ,,,y = mx + b , are you sure slope of x + y = 11 is 1 ?

OpenStudy (anonymous):

did you get my explanation ?

OpenStudy (anonymous):

I did Thanks :)